3% How many caiers ane there in the famiy b. 1 e. 7 7t an affected man has all afected daughters but no affected son t s ain d. Xinked recessive trait 38. The complete, nomal genome for an individual human consis of a 23 chromosomes d. 22 autosoma e. 44 autosomal c *X and Y 39 Which statement applies to the linkage maps of chromosomes They …ys account Mendian grets b. They never account for Mendelian genetics c. The genes on far ends of e Recombination of sister chromatids explains trhe Parents with grey bodies, long wings are crossed with those having black bodies vestigial wings. 470 Grey body, long wing 430 back body, vestigial wing Grey body, vestigial wing 45 black body, long wing 40. Use the following data to determine the recombination frequency beios a, 10% (10 centrorgans) b. 20%(20 certrorgans) c. 15% (15 cert mopars) d. 17% (17 certrorgars) e. 25%(25 centrorgars) 41. What are the a. Barr bodies b. Parental types c. Wild types progeny
Answer #1

36. Ques incomplete

37.x linked dominant trait

38. 22 autosomes plus X or Y

39. The closer 2 loci on chromosome the more likely they will come segregate

40. 45+55/45+55+470+430= 1/10

1/10*100=10% or 10 centimorgan

41. Recombinants